a bus travailing at 30m/s(s) applies its breaks and stops in 6.0s. what is the acceleration of the bus?


1
Expert's answer
2021-03-08T08:24:40-0500
"a=\dfrac{v-v_0}{t}=\dfrac{0-30\ \dfrac{m}{s}}{6\ s}=-5\ \dfrac{m}{s^2}."

The sign minus means that the bus decelerates.